write the augmented matrix for the folowing system of equations
x-7y+z=15
y+6z=14
z=11

Answer by Edwin McCravy(6939)

write the augmented matrix for the 
folowing system of equations
x-7y+z=15
y+6z=14
z=11

 Line 'em up like this:

  x - 7y +  z = 15
       y + 6z = 14
            z = 11

 Put in variable terms with 0 coefficients
 and
 Put in 1 coefficients where they 
 are understood

 1x - 7y + 1z = 15
 0x + 1y + 6z = 14
 0x + 0y + 1z = 11

 Erase all the letters

 1  - 7  + 1  = 15
 0  + 1  + 6  = 14
 0  + 0  + 1  = 11

 Put negative signs on the numbers which have
 minus signs before them, and erase the plus
 signs

 1   -7    1  = 15
 0    1    6  = 14
 0    0    1  = 11

Replace the equal signs with a vertical line

 1   -7    1  | 15
 0    1    6  | 14
 0    0    1  | 11

Place brackets around them

æ1   -7    1  | 15ö	
ç0    1    6  | 14÷
è0    0    1  | 11ø

Edwin